# Hamiltonian learning quantum magnets with non-local impurity tomography

### v1

@@ VI Conclusion
Here we presented a machine learning methodology to infer Hamiltonian parameters that leverages the response of a quantum magnet to local impurity perturbations. Our strategy exploits the non-local changes created by localized impurities in a quantum many-body wavefunction, and in particular the dependence of those changes with respect to an externally tunable magnetic field. Our methodology allows extracting both long-range exchange couplings and anisotropic exchange coupling of quantum magnets, effects that often represent a remarkable challenge with conventional methodologies. Our approach offers a powerful method for extracting complex, non-local information from quantum systems, which is often challenging to access through traditional techniques. We showed that this methodology can be carried out in the presence of different noise sources, showing that this approach is robust even in the presence of realistic perturbations. By adapting our methodology to a single particle problem, we showed that this technique enables learning single-particle dispersion, and can be reinterpreted as a machine-learning inspired quasiparticle interference method. Within the single-particle limit, we showed that this technique allows us to extract spin-orbit coupling terms of the electronic structure from a non-polarized observable, an often challenging effect to extract with conventional quasiparticle interference. Our method showcases the powerful synergy between machine learning and quantum many-body methods, in particular, showing how machine learning techniques trained in computational data can be used to learn Hamiltonians from experimental observables. Our results establish a proof of concept to apply this machine-learning methodology to more complex quantum many-body systems, including higher-dimensional quantum magnets and interacting fermionic systems.
@@ Acknowledgments
We acknowledge financial support from the Academy of Finland Projects Nos. 331342, and 349696, InstituteQ, the Jane and Aatos Erkko Foundation, and the Finnish Quantum Flagship. We acknowledge the computational resources provided by the Aalto Science-IT project. We thank Tiago Antão, Adolfo Fumega, Rouven Koch, Peter Liljeroth and Robert Drost for useful discussions.

### v2

@@ VII Conclusion
Here we presented a machine learning methodology to infer Hamiltonian parameters that leverages the response of a quantum magnet to local impurity perturbations. Our strategy exploits the non-local changes created by localized impurities in a quantum many-body wavefunction, and in particular the dependence of those changes with respect to an externally tunable magnetic field. Our methodology allows extracting both long-range exchange couplings and anisotropic exchange coupling of quantum magnets, effects that often represent a remarkable challenge with conventional methodologies. Our approach offers a powerful method for extracting complex, non-local information from quantum systems, which is often challenging to access through traditional techniques. We showed that this methodology can be carried out in the presence of different noise sources, showing that this approach is robust even in the presence of realistic perturbations. By adapting our methodology to a single particle problem, we showed that this technique enables learning single-particle dispersion, and can be reinterpreted as a machine-learning inspired quasiparticle interference method. Within the single-particle limit, we showed that this technique allows us to extract spin-orbit coupling terms of the electronic structure from a non-polarized observable, an often challenging effect to extract with conventional quasiparticle interference.
Our method showcases the powerful synergy between machine learning and quantum many-body methods, in particular, showing how machine learning techniques trained in computational data can be used to learn Hamiltonians from experimental observables. Our results establish a proof of concept to apply this machine-learning methodology to more complex quantum many-body systems, including higher-dimensional quantum magnets and interacting fermionic systems.
Data Availability
The datasets and trained models are publicly available at Ref. [104].
@@ Acknowledgments
We acknowledge financial support from the Research Council of Finland Projects Nos. 331342, and 358088, InstituteQ, the Jane and Aatos Erkko Foundation, and the Finnish Quantum Flagship. We acknowledge the computational resources provided by the Aalto Science-IT project. We thank Tiago Antão, Adolfo Fumega, Rouven Koch, Peter Liljeroth and Robert Drost for useful discussions.
Appendix A Neural Network Training
A.1 Fermionic Chain
For the fermionic chain, we employed a fully connected NN architecture. The physical observable used as input to train the network is the spectral function, as defined in Section III.1. We computed the spectral function using a Green’s function formalism [105], where the energy smearing is taken as $\Delta=1\times 10^{-2}t$. We generated a dataset of 5000 samples, with 3000 samples designated for training and 1000 samples for testing. The training process spanned 200 epochs, utilizing a batch size of 32. The Adam optimizer was used for optimization, with a learning rate set to $10^{-4}$. The detailed layer structure of the NN is summarized in Table 1.
Layer Neurons Activation
Input 900 -
Dense 512 ReLU
Dense 512 ReLU
Dense 256 ReLU
Dense 256 ReLU
Dense 128 ReLU
Dense 128 ReLU
Dense 64 ReLU
Dense 64 ReLU
Dense 32 ReLU
Dense 32 ReLU
Dense 16 ReLU
Dense 16 ReLU
Output 3 Linear
Table 1: Architecture of the NN used for learning the parameters of the fermionic system.
The input data for the NN was prepared by considering a chain of 150 sites. We evaluated the spectral function for three different voltage biases, $V\in[-0.5,0,0.5]$, and then applied a Fourier transform to the spectral function data. This preprocessing step resulted in an input size of 900 for the first layer of the NN. The loss function used for training was the mean squared error (MSE). The chosen parameter ranges are as follows: $\mu$ and $\lambda_{R}$ ranged from $[0,1]$, while $t_{2}$ spanned $[0,2]$.
The training process followed a two-step approach. Initially, the model was trained solely on noiseless data, employing an early stopping criterion based on validation loss within the first 5 epochs. After this phase, magnetic noise was added to the original data set at levels of $\Delta^{2}_{B}=0.05$ and $\Delta^{2}_{B}=0.1$. We then resumed training the same NN for an additional 200 epochs. This methodology allows us to assess the robustness of the model under varying noise levels. At the end of the process, we obtained three models: one trained exclusively on noiseless data and two additional models fine-tuned with increasing levels of noise. The workflow of our NN training for the fermionic chain is illustrated in Fig. 4a.
A.2 Spin Chain
For the spin chain system, we adopted the same fully connected NN architecture, utilizing the static spin correlator as input, as defined in Section III.2. The dataset consisted of 2000 samples, with 1200 samples for training and 400 samples for testing. The training process lasted 100 epochs, using the Adam optimizer with a learning rate of $10^{-4}$ and batch size of 32. The detailed layer structure of the NN is summarized in Table 2.
Layer Neurons Activation
Input 20 -
Dense 32 ReLU
Dense 32 ReLU
Dense 16 ReLU
Dense 16 ReLU
Dense 8 ReLU
Dense 8 ReLU
Output 3 Linear
Table 2: Architecture of the NN used for learning the parameters of the spin chain system.
The input data was derived from a spin chain of 21 sites. We evaluated the static spin correlator for 12 different external magnetic fields in the range $B\in[0,5]$, resulting in an initial input size of 252. PCA was then applied to reduce the input size to 20 components. The loss function used for training was MSE, with the parameter ranges for $J_{2}$ and $J_{z}$ spanning $[-0.5,0.5]$, and $J_{DM}$ spanning $[0,1]$.
We adopted the same two-step training strategy as in the fermionic chain case. Initially, the model was trained on noiseless data, with early stopping applied within the first 5 epochs based on validation loss. Subsequently, magnetic noise at levels of $\Delta^{2}_{B}=0.05$ and $\Delta^{2}_{B}=0.1$ was added to the dataset. PCA was reapplied, and training resumed for an additional 100 epochs. This process allowed us to systematically evaluate the model’s robustness against noise. The workflow of the NN training for the spin chain is illustrated in Fig. 6a.
We also explored an alternative network architecture with separate branches to learn $J_{2}$, $J_{z}$, and $J_{DM}$. However, this approach did not produce significant performance improvements. Therefore, we adopt a simultaneous learning strategy that ensures consistency across all parameters while maintaining comparable or superior performance.